If the diffusion coefficient of oxygen in the vitreous humour in the eye is 4.2×10−7 m2 s−1 how long does it take oxygen to diffuse an average distance of 1.0 mm through the vitreous humour? (to 2 s.f and in s)
1
Expert's answer
2020-04-24T13:35:50-0400

D=4.2×10-7 m2 s-1

d=1mm

mass m of oxygen=32g

"\delta m=\frac {D*\delta t}{d}"

dt=(32*1)/4.2×10-7

dt=7.6*107